OpenStudy (anonymous):

Solve for x: 3x - 5 = 2x + 6.

OpenStudy (anonymous):

\(\Large 3x - 5 = 2x + 6\) \(\Large 3x-2x=6+5\) \(\Large x=11\)

OpenStudy (anonymous):

3x - 2x = 6 + 5

OpenStudy (anonymous):

just rearrange the terms a bit

OpenStudy (anonymous):

6 + 5 = 11 3 * 11 = 33 2 * 11 = 22 33 - 22 = 11

OpenStudy (anonymous):

so x = 11